satellite73 (satellite73):

what is the smallest positive integer \(n\) such that out of the \(n\) unit fractions \(\frac{1}{k}\) where \(1\le k \le n\) exactly half the fractions give a terminating decimal/

OpenStudy (anonymous):

someone posted this, not sure it is easy or hard i am up to 6

myininaya (myininaya):

so if n=2 we have two fractions 1/1,1/2 Both of these terminate so that is more than half that give a terminating decimal

myininaya (myininaya):

So based on my response this means I have the right interpretation?

OpenStudy (anonymous):

ooho i think i got it!

myininaya (myininaya):

Not the correct answer just correct interpretation.. I would think n would have to be even then.

OpenStudy (anonymous):

i guess so i copied the question verbatim

myininaya (myininaya):

n=2i \[\frac{1}{1} ,\frac{1}{2}, \frac{1}{3}, \cdots, \frac{1}{2i-1}, \frac{1}{2i}\] so we want half of these to terminate for some integer i

OpenStudy (anonymous):

i made two lists until i got half and half 1, 1/2 1/4 1/5 1/8 1/10 1/3 1/6 1/7 1/9 1/11 1/12

OpenStudy (anonymous):

too bad the asker has flown the coop

myininaya (myininaya):

well look at you you get a brownie and you can have one of my kitties

OpenStudy (anonymous):

i have never kittled!

zepdrix (zepdrix):

Hmm, neat problem :)

myininaya (myininaya):

you should kittled! Kittling is the cutest.

myininaya (myininaya):

I wonder if this was a number theory question or if it was a question that a math teacher came up with to test who knows what a terminating decimal is.

OpenStudy (anonymous):

probably the latter i have another one that i will post now it is slightly more difficult but only slightly

myininaya (myininaya):

I guess they would have to be able to read that which they probably wouldn't be able to.